if a 55 n force is required to move an object that weighs 93 n across a level surface at a constant speed, what is the coefficient of kinetic friction
1
Expert's answer
2019-02-06T13:34:30-0500

By the definition of the coefficient of kinetic friction we have:

"\mu_k = \frac{F_{appl}}{W},"

here,

"F_{appl}"

is the horizontal pushing force that required to move an object across a level surface at a constant speed,

"W"

is the weight of the object directed downward.

Then, we get:

"\mu_k = \frac{F_{appl}}{W} = \frac{55 N}{93 N} = 0.59."

Answer:

"\mu_k = 0.59"

.